OpenStudy (anonymous):

Indefinite integral wrt x of x sqrt(1+x)?

OpenStudy (anonymous):

2/5*(x + 1)^(5/2) - 2/3*(x + 1)^(3/2) + C substitute u = 1 + x

myininaya (myininaya):

\[\int\limits_{}^{}x*\sqrt{x+1} dx\] let u=x+1 du=1 dx if u=x+1, then u-1=x \[\int\limits_{}^{}(u-1) \sqrt{u} du\] =\[\int\limits_{}^{}(u^{1+\frac{1}{2}}-u^\frac{1}{2} )du\] \[\int\limits_{}^{}(u^\frac{3}{2}-u^\frac{1}{2})du\]